OpenStudy (anonymous):

Two cars start moving from the same point. One travels south at 60 miles/hr and the other travels west at 25 miles/hr. At what rate is the car distance between the cars increasing 2 hours later. No calculators, no cheating!! Right answer gets a cookie! Go!

OpenStudy (jack1):

no fair, wio's part calculator, he's like sheldon from BB theory ;D

OpenStudy (anonymous):

if wio gets the right answer he'd dq'd

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

\[ z^2=x^2+y^2 \]So \[ 2zz'=2xx'+2yy' \]Dividing by \(2\)\[ zz'=25x+60y\implies z'=\frac{xx'+yy'}{\sqrt{x^2+y^2}{}}=\frac{25x+60y}{\sqrt{x^2+y^2}{}} \]

OpenStudy (anonymous):

We can use \(x=25\times 2=50\) and \(y=60\times 2=120\).

OpenStudy (anonymous):

It's not going to be a rational number.

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

\(2500+14400=16900 = 130^2\)?

OpenStudy (anonymous):

Draw a picture wio!

OpenStudy (anonymous):

Ok

OpenStudy (timaashorty):

WIO ! WAY TO GO! I'LL BE CHEERING YOU ON ! YOU COULD DO THIS !! I WISH I COULD GIVE YOU 10 MEDALS! CUZ I WOULD. WIO SWEETIE YOU WILL GET THE ANSWER!

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (timaashorty):

WIO! WIO! WIO! WIO! BEST HELPER IN HERE! WIO WIO. OH MY WIO. WIO I LOVE YOU.

OpenStudy (timaashorty):

WIO YOU ARE SIMPLY. AWESOME.

OpenStudy (primeralph):



OpenStudy (timaashorty):

PRIME BACK OFF. WIO GOT THIS.

OpenStudy (shamil98):

L is here have no fear!

OpenStudy (anonymous):

\[ z'=\frac{1250+7200}{130}=\frac{845}{13} \]

OpenStudy (anonymous):

Wio wins!

OpenStudy (anonymous):

\[ z'=65 \]

OpenStudy (timaashorty):

WIO, YOUR BABYGIRL HIS HERE. YOU COULD DO IT

OpenStudy (timaashorty):

*IS

OpenStudy (shamil98):

Let variable x be the distance car A travels in t hours, and variable y the distance car B travels in t hours. Let variable L be the distance between cars A and B after t hours.

OpenStudy (timaashorty):

GET OUT SHAM! MY BOO WIO DID IT !

OpenStudy (shamil98):

60 - 25 = 35 so the equation for L would be L = √x^2 + (35-y)^2

OpenStudy (anonymous):

Well done @wio ^.^

OpenStudy (timaashorty):

GOOOOOOOOOOOOOOOOOOOALLL ! WIO ACTUALLY DID IT! EVERYONE SHOULD GIVE HIM A MEDAL!

OpenStudy (luigi0210):

Poor @wio

OpenStudy (timaashorty):

EVERY. SINGLE. PERSON SEEING THIS. NEED. TO. GIVE HIM. A. MEDAL.

OpenStudy (primeralph):

Other people need help. Get to it.

OpenStudy (shamil98):

well since the rate is after two hours it would be the same as x = 50t and y = 120t

OpenStudy (timaashorty):

PRIME YOU DID AMAZING TOO. CONGRATS!

OpenStudy (primeralph):

@timaashorty Are you new here?

OpenStudy (anonymous):

hahahahha prime

OpenStudy (timaashorty):

NOPE (:

OpenStudy (shamil98):

L = √x^2 + ( 70 - y)^2